Please help me solve this question.......The directions are asking you to write yes or no to show whether each proportion is a true proportion. Here is the question.........


6/13 = 18/39

Answer by Alan3354(69443)

6/13 = 18/39
------------
Hint: 6*3 = 18 and 13*3 = 39
(6/13)*(3/3) = ?